# Operator Learning for Continuous Spatial-Temporal Model with Gradient-Based and Derivative-Free Optimization Methods

### v2

@@ Abstract
Partial differential equations are often used in the spatial-temporal modeling of complex dynamical systems in many engineering applications. In this work, we build on the recent progress of operator learning and present a data-driven modeling framework that is continuous in both space and time. A key feature of the proposed model is the resolution-invariance with respect to both spatial and temporal discretizations, without demanding abundant training data in different temporal resolutions. To improve the long-term performance of the calibrated model, we further propose a hybrid optimization scheme that leverages both gradient-based and derivative-free optimization methods and efficiently trains on both short-term time series and long-term statistics. We investigate the performance of the spatial-temporal continuous learning framework with three numerical examples, including the viscous Burgers’ equation, the Navier–Stokes equations, and the Kuramoto–Sivashinsky equation. The results confirm the resolution-invariance of the proposed modeling framework and also demonstrate stable long-term simulations with only short-term time series data. In addition, we show that the proposed model can better predict long-term statistics via the hybrid optimization scheme with a combined use of short-term and long-term data.
@@
Operator learning, Spatial-temporal model, Gradient-based optimization, Ensemble Kalman methods
@@ 1 Introduction
Complex dynamical systems are ubiquitous in many science and engineering applications, e.g., mixing phenomena in atmospheric and ocean dynamics gill1982atmosphere ; fernando1991turbulent ; vallis2017atmospheric , physical and chemical processes of energy conversion dimotakis2005turbulent , and drones operating in turbulent flows shi2019neural . For most of these systems, numerically simulating the governing equations derived by first principles is still infeasible in the foreseeable future, and thus closure models are inevitably needed to account for those unresolved degrees of freedom. Many of the classical closure models are calibrated with simplified settings and are known to have limited predictive capability, mainly due to the lack of enough expressive power in the model form and the empirical calibration to a small amount of data. In the past few decades, we have witnessed the rapid growth of high-fidelity simulation and experimental data and the great advance of machine learning methods, which motivated the development of data-driven modeling techniques brunton2016discovering ; kutz2016dynamic ; wang2017physics ; wu2018physics ; duraisamy2019turbulence ; brunton2020machine ; schneider2021learning ; gupta2021neural ; brunton2022data ; chen2023causality ; chen2023ceboosting with the aim of improving or even replacing the classical models by neural-network-based models.
@@
Standard machine learning methods have achieved great success in the past decade, e.g., convolutional neural networks lecun1989handwritten ; lecun1998gradient for tasks such as computer vision (CV), and recurrent neural networks funahashi1993approximation ; hochreiter1997long ; jaeger2001echo ; maass2002real for tasks such as natural language processing (NLP). More recently, transformer-based models have demonstrated even better performance in both CV and NLP tasks. Inspired by the success of these methods in standard machine learning tasks, many previous works explored using these methods to improve the modeling of dynamical systems. Although preliminary success has been promisingly demonstrated, these standard machine learning methods often assume fixed discretization in both space and time, which turns out to be a strong limitation for the modeling of complex dynamical systems. The main reason for such a limitation is two-fold: (i) the available data sources are unlikely in a consistent resolution, and (ii) the efficient characterization of the system state usually demands a non-uniform resolution. For instance, the data sources of earth system include high-fidelity simulations wedi2020baseline with a fine spatial resolution such as $O(100\textrm{m})$ to $O(1\textrm{km})$, satellite images across a wide range of resolutions with the finest one of $O(10\textrm{m})$, and the data collected by observatories that sparsely distributed across the earth. The temporal resolutions as well vary among high-fidelity simulations, satellite images, and observatory data. This multiscale nature of different data sources is also present in many engineering applications, e.g., wind farms with both LiDAR data and high-fidelity simulations of atmospheric boundary layer porte2011large ; lu2011large . On the other hand, the system states of complex dynamical often have multiscale features intrinsically, for which an adaptive resolution is often the more efficient way of characterizing the system state than a uniform resolution.
@@
The multiscale nature of both the system state and the available data sources of complex dynamical systems has motivated the development of continuous form machine learning methods for the modeling and simulation of those systems. For instance, Fourier neural operator (FNO) li2021fourier employs Fourier and inverse Fourier transformations to construct Fourier layers, which are used as building blocks to approximate an operator (i.e., a mapping between Banach spaces) in the continuous form. Essentially, FNO belongs to a general category of neural operators kovachki2023neural , i.e., using neural networks to approximate the integral operator as the basic component of the approximation of a more general operator. Other operator learning methods that belong to this category are graph neural operators (GNO) anandkumar2020neural ; li2020multipole and low-rank neural operators (LNO), which connects to another popular operator learning framework known as DeepONet lu2019deeponet ; lu2021learning . More specifically, DeepONet introduced branch net and trunk net to implement an approximation of operator based on the universal approximation theorem of operators. There are also other recent works that construct neural-network-based operators via a wavelet-based approximation of integral operators gupta2021multiwavelet or reconstruction from other types of sub-networks patel2022variationally . Comparisons between operator learning approaches have been studied in lu2022comprehensive ; kovachki2023neural . The approximation error has been studied in cao2023residual with a residual-based error correction method being proposed. Physics constraints li2021physics ; wang2021learning and derivative information o2022derivate were demonstrated as additional information sources that can enhance the performance of operator learning methods. There have been many other interesting extensions of the standard operator learning frameworks, e.g., solving problems on general geometries li2022fourier , learning of nonlinear non-autonomous dynamical systems lin2023learning , and improving operator learning techniques with the recent success of large language models yang2023context ; liu2023prose , to name a few. Although these methods can be used to characterize the solution operator of a dynamical system, it only work with a fixed and uniform temporal resolution if standard operator learning methods are used. By taking the prediction lead time as an additional input, the operator learning methods can potentially work with non-uniform temporal resolution but often demand abundant data with different temporal resolutions for the training.
@@
On the other hand, neural ordinary differential equation (ODE) chen2018neural provided a temporally continuous framework of machine learning methods for the modeling and simulation of dynamical systems. Recently, the combined use of neural ODE and neural operator has been explored by cho2022neural in the context of standard machine learning tasks such as classification and computer vision. In terms of spatial-temporal modeling of dynamical systems, the key relevant concepts of neural ODE are: (i) neural networks can be used to characterize the unknown vector field of a finite-dimensional dynamical system, and (ii) the unknown coefficients of the neural network can be trained via gradient descent using either the adjoint method or automatic differentiation. The idea of modeling continuous dynamics via machine learning methods was also discussed in weinan2017 . In the past few years, several works maulik2020time ; portwood2019turbulence ; linot2022data ; djeumou2022neural have explored the use of neural ODE in modeling dynamical systems and demonstrated promising results. Although neural ODE provides the flexibility of handling data with a non-uniform resolution in time, it has been shown in linot2023stabilized that the use of a standard network in neural ODE can lead to long-term instability, mainly due to the amplification of high wavenumber components, when applying the trained network and simulating the modeled dynamical system. To address the long-term stability, the unknown vector field was modeled by linear and nonlinear parts with standard neural networks in linot2023stabilized and the training via neural ODE demonstrated more stable long-term simulations. In this work, we show that the stable long-term simulations of neural ODE models can alternatively be achieved by filtering out some high wavenumber components through the construction of a neural dynamical operator, for which the Fourier neural operator li2021fourier serves the purpose nicely.
@@
Although neural ODE provides an efficient tool for training a model to match the short-term behavior of an unknown dynamical system, training the model to also quantitatively match the long-term system behavior still has some challenges. One challenge is on the computational side, that the long-term training would require huge memory costs paszke2017automatic ; baydin2018automatic ; margossian2019review via backpropagation or potential stability issues in the backward-in-time simulation via the adjoint method. Another challenge is on the problem formulation side, that the long-term system behavior tends to be more sensitive to the small changes in the model, especially for chaotic systems, making the standard adjoint method sometimes even infeasible wang2014least . These challenges motivate us to explore an alternative optimization approach that is both efficient and robust for matching the long-term behaviors between the model and the true dynamical system. More specifically, we choose the ensemble Kalman inversion (EKI) method that was proposed in iglesias2013ensemble . Unlike the backpropagation or the adjoint method that is designed for the gradient descent optimization, ensemble Kalman inversion is derivative-free, parallelizable, and robust with noises of the data and chaos or uncertainties of the system schneider2021learning ; wu2023learning . In the past decade, many developments have been made to enhance Kalman inversion methods, both in theory schillings2017analysis ; ding2021ensemble ; calvello2022ensemble and in algorithms, such as various types of regularizations (e.g., linear equalities and inequalities albers2019ensemble , Tikhonov chada2020tikhonov , sparsity schneider2022ensemble , and $\ell_{p}$ lee2021l_p , and known physics zhang2020regularized ), uncertainty quantification with Langevin dynamics for sampling garbuno2020interacting , and using other types of Kalman filters huang2022iterated . More recently, the ensemble Kalman inversion was also explored in bottcher2023gradient for the training of neural ODE. Instead of merely using EKI to train a neural ODE, we investigate a hybrid optimization approach that uses standard gradient-based method for the short-term trajectory matching and the EKI method for long-term statistics matching.
@@
In summary, we develop a spatially and temporally continuous framework for the data-driven modeling of dynamical systems with spatial fields as system states, by leveraging the success of operator learning and neural ODE. The framework is tested with three chaotic systems, including 1-D Burgers’ equation, 2-D Navier–Stokes equation, and Kuramoto–Sivashinsky equation. The first two examples focus on the performances of neural dynamical operator trained with short-term data, to demonstrate the merits of the neural dynamical operator in terms of (i) spatial-temporal resolution-invariant and (ii) stable long-term simulations. The third example demonstrates the merit of an hybrid optimization method that leverages both short-term and long-term data. The key contributions of this work are summarized below:
@@ 2.1 Problem Setting
with spatial variable $\bm{x}\in D_{{\bm{x}}}\subseteq\mathbb{R}^{d_{\bm{x}}}$, temporal variable $t\in[0,T]\subset\mathbb{R}$, state function $\bm{u}(\cdot,\cdot)\in\mathcal{U}(D_{\bm{x}}\times[0,T];\mathbb{R}^{d_{\bm{u}}})$, spatial profile of state function at time $t$ is $\bm{u}(\cdot,t)\in\mathcal{U}_{t}(D_{\bm{x}};\mathbb{R}^{d_{\bm{u}}})$, and $\mathcal{G}:\mathcal{U}_{t}\times[0,T]\mapsto\mathcal{U}_{t}$ is a non-linear operator. If the system is autonomous, i.e., the system does not depend on time, $\mathcal{G}$ would become a non-linear spatial operator. $\mathbb{R}^{d}$ is a real vector space with $d$ dimension, $D_{\bm{x}}$ is a bounded domain, $\mathcal{U}$ and $\mathcal{U}_{t}$ are separable Banach spaces of function taking values in $\mathbb{R}^{d_{\bm{u}}}$.
@@
Assuming that the operator $\mathcal{G}$ in Eq. (2.1) is unknown, we propose to learn a data-driven dynamical operator $\mathcal{\tilde{G}}$ that approximates $\mathcal{G}$, based on the time series data of $\bm{u}$ and a combined use of neural ODE chen2018neural and neural operator li2021fourier . To enhance the performance of the learned model, the data-driven operator is trained to match both short-term trajectories and long-term statistics from the true system. A hybrid scheme of gradient-based and derivative-free methods is then demonstrated to facilitate efficient training with the combined types of data.
@@
Given a time series $\{\bm{u}_{t_{n}}\}_{n=0}^{N}$ of system state $\bm{u}$ in Eq. (2.1), where $\bm{u}_{t_{n}}$ serves as a short form of $\bm{u}(\cdot,t_{n})$, we aim to build a continuous spatial-temporal model for the system by constructing a parametric map $\mathcal{\tilde{G}}$ with parameters $\theta$ to approximate the non-linear operator $\mathcal{G}$. The continuous spatial-temporal model would then be:
$\frac{\partial\tilde{\bm{u}}(\bm{x},t)}{\partial t}=\mathcal{\tilde{G}}(\tilde{\bm{u}}(\bm{x},t),t;\theta).$ (2.2)
To match with the set of $N$ time series, the $\tilde{\mathcal{G}}$ can be obtained by solving an optimization problem:
$\min_{\theta}\sum_{n=0}^{N}L(\bm{u}_{t_{n}},\ \tilde{\bm{u}}_{t_{n}}).$ (2.3)
@@
where $\tilde{\bm{u}}_{t_{n}}=\bm{u}_{t_{0}}+\int_{t_{0}}^{t_{n}}\tilde{\mathcal{G}}(\bm{u},t;\theta)\mathop{}\!\mathrm{d}t$ is the predicted states, $\bm{u}_{t_{n}}$ is the observed system states from the true system, and $L:\mathcal{U}_{t}\times\mathcal{U}_{t}\mapsto\mathbb{R}$ is a loss functional. In practice, we may have a long time series of states as data for a dynamical system. Instead of directly matching the long time series by solving the optimization problem in Eq. (2.3), we divide the data into batches of short time series, with which we can use mini-batch training to obtain the parameters $\theta$.
@@
To capture a long-term statistical property of the true dynamical system, $\tilde{\mathcal{G}}$ can be calibrated by solving another optimization problem:
$\min_{\theta}L_{l}(\beta(\{\bm{u}_{t_{n}}\}_{n=0}^{l}),\ \beta(\{\tilde{\bm{u}}_{t_{n}}\}_{n=0}^{l})).$ (2.4)
where $\beta:\mathcal{U}^{l+1}\mapsto\mathbb{R}^{d_{\beta}}$ is a long-term statistics functional of the true system and the modeled one, and $L_{l}:\mathbb{R}^{d_{\beta}}\times\mathbb{R}^{d_{\beta}}\mapsto\mathbb{R}$ is a loss function. Instead of taking the probability measure of the Banach space $\mathcal{U}$ as its input, the functional $\beta$ only estimates the long-term statistics of $\bm{u}$ and $\tilde{\bm{u}}$ from the time series data $\{\bm{u}_{t_{n}}\}_{n=0}^{l}$ and $\{\tilde{\bm{u}}_{t_{n}}\}_{n=0}^{l}$, with $l\gg N$ in Eq. (2.3).

### v3

# Neural Dynamical Operator: Continuous Spatial-Temporal Model with Gradient-Based and Derivative-Free Optimization Methods

@@ Abstract
Data-driven modeling techniques have been explored in the spatial-temporal modeling of complex dynamical systems for many engineering applications. However, a systematic approach is still lacking to leverage the information from different types of data, e.g., with different spatial and temporal resolutions, and the combined use of short-term trajectories and long-term statistics. In this work, we build on the recent progress of neural operator and present a data-driven modeling framework called neural dynamical operator that is continuous in both space and time. A key feature of the neural dynamical operator is the resolution-invariance with respect to both spatial and temporal discretizations, without demanding abundant training data in different temporal resolutions. To improve the long-term performance of the calibrated model, we further propose a hybrid optimization scheme that leverages both gradient-based and derivative-free optimization methods and efficiently trains on both short-term time series and long-term statistics. We investigate the performance of the neural dynamical operator with three numerical examples, including the viscous Burgers’ equation, the Navier–Stokes equations, and the Kuramoto–Sivashinsky equation. The results confirm the resolution-invariance of the proposed modeling framework and also demonstrate stable long-term simulations with only short-term time series data. In addition, we show that the proposed model can better predict long-term statistics via the hybrid optimization scheme with a combined use of short-term and long-term data.
@@
Complex dynamical system , Neural operator , Spatial-temporal model , Gradient-based optimization , Ensemble Kalman methods
@@ 1 Introduction
Complex dynamical systems are ubiquitous in many science and engineering applications, e.g., mixing phenomena in atmospheric and ocean dynamics [1, 2, 3], physical and chemical processes of energy conversion [4], and drones operating in turbulent flows [5]. For most of these systems, numerically simulating the governing equations derived by first principles is still infeasible in the foreseeable future, and thus closure models are inevitably needed to account for those unresolved degrees of freedom. Many of the classical closure models are calibrated with simplified settings and are known to have limited predictive capability, mainly due to the lack of enough expressive power in the model form and the empirical calibration to a small amount of data. In the past few decades, we have witnessed the rapid growth of high-fidelity simulation and experimental data and the great advance of machine learning methods, which motivated the development of data-driven modeling techniques [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] with the aim of improving or even replacing the classical models by neural-network-based models.
@@
Standard machine learning methods have achieved great success in the past decade, e.g., convolutional neural networks [19, 20] for tasks such as computer vision (CV), and recurrent neural networks [21, 22, 23, 24] for tasks such as natural language processing (NLP). More recently, transformer-based models have demonstrated even better performance in both CV and NLP tasks. Inspired by the success of these methods in standard machine learning tasks, many previous works explored using these methods to improve the modeling of dynamical systems. Despite the initial promising results, standard machine learning methods often assume fixed discretization in both space and time, which presents a significant constraint for accurately modeling complex dynamical systems. The main reason for such a limitation is two-fold: (i) the available data sources are unlikely in a consistent resolution, and (ii) the efficient characterization of the system state usually demands a non-uniform resolution. For instance, the data sources of earth system include high-fidelity simulations [25] with a fine spatial resolution such as $O(100\textrm{m})$ to $O(1\textrm{km})$, satellite images across a wide range of resolutions with the finest one of $O(10\textrm{m})$, and the data collected by observatories that sparsely distributed across the earth. The temporal resolutions also vary among high-fidelity simulations, satellite images, and observatory data. This multiscale nature of different data sources is also present in many engineering applications, e.g., wind farms with both LiDAR data and high-fidelity simulations of atmospheric boundary layer [26, 27]. On the other hand, the system states of complex dynamics often have multiscale features intrinsically, for which an adaptive resolution is often the more efficient way of characterizing the system state than a uniform resolution.
@@
The multiscale nature of both the system state and the available data sources of complex dynamical systems has motivated the development of continuous form machine learning methods for the modeling and simulation of those systems. For instance, Fourier neural operator (FNO) [28] employs Fourier and inverse Fourier transformations to construct Fourier layers, which are used as building blocks to approximate an operator (i.e., a mapping between Banach spaces) in the continuous form. Essentially, FNO belongs to a general category of neural operators [29], i.e., using neural networks to approximate the integral operator as the basic component of the approximation of a more general operator. Other neural operator methods that belong to this category are graph neural operators (GNO) [30, 31] and low-rank neural operators (LNO), which connects to another popular neural operator framework known as DeepONet [32, 33]. More specifically, DeepONet introduced branch net and trunk net to implement an approximation of operator based on the universal approximation theorem of operators. There are also other recent works that construct neural-network-based operators via a wavelet-based approximation of integral operators [34] or reconstruction from other types of sub-networks [35]. Comparisons between neural operator approaches have been studied in [36, 29]. The approximation error has been studied in [37] with a residual-based error correction method being proposed. Physics constraints [38, 39] and derivative information [40] were demonstrated as additional information sources that can enhance the performance of neural operator methods. There have been many other interesting extensions of the standard neural operator frameworks, e.g., solving problems on general geometries [41], learning of nonlinear non-autonomous dynamical systems [42], and improving neural operator techniques with the recent success of large language models [43, 44], to name a few. Although these methods can be used to characterize the solution operator of a dynamical system, they only work with a fixed and uniform temporal resolution if standard neural operator methods are used. By taking the prediction lead time as an additional input, the neural operator methods can potentially work with non-uniform temporal resolution but often demand abundant data with different temporal resolutions for the training.
@@
On the other hand, neural ordinary differential equation (ODE) [45] provided a temporally continuous framework of machine learning methods for the modeling and simulation of dynamical systems. Recently, the combined use of neural ODE and neural operator has been explored by [46] in the context of standard machine learning tasks such as classification and computer vision. In terms of spatial-temporal modeling of dynamical systems, the key relevant concepts of neural ODE are: (i) neural networks can be used to characterize the unknown vector field of a finite-dimensional dynamical system, and (ii) the unknown coefficients of the neural network can be trained via gradient descent using either the adjoint method or automatic differentiation. The idea of modeling continuous dynamics via machine learning methods was also discussed in [47]. In the past few years, several works [48, 49, 50, 51] have explored the use of neural ODE in modeling dynamical systems and demonstrated promising results. Although neural ODE provides the flexibility of handling data with a non-uniform resolution in time, it has been shown in [52] that the use of a standard network in neural ODE can lead to long-term instability, mainly due to the amplification of high wavenumber components, when applying the trained network and simulating the modeled dynamical system. To address the long-term stability, the unknown vector field was modeled by linear and nonlinear parts with standard neural networks in [52] and the training via neural ODE demonstrated more stable long-term simulations. In this work, we show that the stable long-term simulations of neural ODE models can alternatively be achieved by filtering out some high wavenumber components through the construction of a neural dynamical operator, for which the Fourier neural operator [28] serves the purpose nicely.
@@
Although neural ODE provides an efficient tool for training a model to match the short-term behavior of an unknown dynamical system, training the model to also quantitatively match the long-term system behavior still has some challenges. One challenge is on the computational side, that the long-term training would require huge memory costs [53, 54, 55] via backpropagation or potential stability issues in the backward-in-time simulation via the adjoint method. Another challenge is on the problem formulation side, that the long-term system behavior tends to be more sensitive to the small changes in the model, especially for chaotic systems or systems with discontinuities, making the standard adjoint method sometimes even infeasible [56]. These challenges motivate us to explore an alternative optimization approach that is both efficient and robust for matching the long-term behaviors between the model and the true dynamical system. More specifically, we choose the ensemble Kalman inversion (EKI) method that was proposed in [57]. Unlike the backpropagation or the adjoint method that is designed for the gradient descent optimization, ensemble Kalman inversion is derivative-free, parallelizable, and robust with noises of the data and chaos or uncertainties of the system [12, 58]. In the past decade, many developments have been made to enhance Kalman inversion methods, both in theory [59, 60, 61] and in algorithms, such as various types of regularizations (e.g., linear equalities and inequalities [62], Tikhonov [63], sparsity [64], and $\ell_{p}$ [65], and known physics [66]), uncertainty quantification with Langevin dynamics for sampling [67], and using other types of Kalman filters [68]. More recently, the ensemble Kalman inversion was also explored in [69] for the training of neural ODE. Instead of merely using EKI to train a neural ODE, we investigate a hybrid optimization approach that uses the standard gradient-based method for short-term trajectory matching and the EKI method for long-term statistics matching.
@@
In summary, we develop a spatially and temporally continuous framework called neural dynamical operator for the data-driven modeling of dynamical systems with spatial fields as system states, by leveraging the success of neural operator and neural ODE. Compared with classical methods, the proposed framework is more capable of capturing the underlying dynamics from data for both short-term state prediction and long-term statistics matching with resolution-invariant feature. The framework is tested with three dynamical systems, including 1-D Burgers’ equation, 2-D Navier–Stokes equation, and Kuramoto–Sivashinsky equation. The first two examples focus on the performances of neural dynamical operator trained with short-term data, to demonstrate the merits of the neural dynamical operator in terms of (i) spatial-temporal resolution-invariant and (ii) stable long-term simulations. The third example demonstrates the merit of a hybrid optimization method that leverages both short-term and long-term data to calibrate the neural dynamical operator. The key contributions of this work are summarized below:
@@ 2.1 Problem Setting
with spatial variable $\bm{x}\in D_{{\bm{x}}}\subseteq\mathbb{R}^{d_{\bm{x}}}$, temporal variable $t\in[0,T]\subset\mathbb{R}$, state function $\bm{u}(\cdot,\cdot)\in\mathcal{U}(D_{\bm{x}}\times[0,T];\mathbb{R}^{d_{\bm{u}}})$, spatial profile of state function at time $t$ is $\bm{u}(\cdot,t)\in\mathcal{U}_{t}(D_{\bm{x}};\mathbb{R}^{d_{\bm{u}}})$, and $\mathcal{G}:\mathcal{U}_{t}\times[0,T]\mapsto\mathcal{U}_{t}$ is a non-linear operator. If the system is autonomous, i.e., the system does not depend on time, $\mathcal{G}$ would become a non-linear spatial operator. $\mathbb{R}^{d}$ is a real vector space with $d$ dimension, $D_{\bm{x}}$ is a bounded domain, $\mathcal{U}$ and $\mathcal{U}_{t}$ are separable Banach spaces of functions taking values in $\mathbb{R}^{d_{\bm{u}}}$.
@@
Assuming that the operator $\mathcal{G}$ in Eq. (2.1) is unknown, we propose to learn a data-driven dynamical operator $\mathcal{\tilde{G}}$ that approximates $\mathcal{G}$, based on the time series data of $\bm{u}$ and a combined use of neural ODE [45] and neural operator [28]. To enhance the performance of the learned model, the data-driven operator is trained to match both short-term trajectories and long-term statistics from the true system. A hybrid scheme of gradient-based and derivative-free methods is then demonstrated to facilitate efficient training with the combined types of data.
@@
Given a time series $\{\bm{u}(t_{n})\}_{n=0}^{N}$ of system state $\bm{u}$ in Eq. (2.1), where $\bm{u}(t):=\bm{u}(\cdot,t)|D_{\bm{x}}^{(M)}\in\mathbb{R}^{d_{\bm{u}}\times M}$ is discretization of $\bm{u}(\cdot,t)$ with $D_{\bm{x}}^{(M)}=\{x_{1},...,x_{M}\}\subset D_{\bm{x}}$ is a $M$-point discretization of the domain $D_{\bm{x}}$, we aim to build a continuous spatial-temporal model for the system by constructing a parametric map $\mathcal{\tilde{G}}$ with parameters $\bm{\theta}$ to approximate the non-linear operator $\mathcal{G}$. The continuous spatial-temporal model would then be:
$\frac{\partial\tilde{\bm{u}}(\bm{x},t)}{\partial t}=\mathcal{\tilde{G}}(\tilde{\bm{u}}(\bm{x},t),t;\bm{\theta}).$ (2.2)
To match with the time series data $\{\bm{u}(t_{n})\}_{n=0}^{N}$, the unknown parameters $\bm{\theta}$ in $\tilde{\mathcal{G}}$ can be calibrated by minimizing the short-term loss for trajectory matching:
$L_{s}:=\sum_{n=0}^{N_{s}}J_{s}(\bm{u}(t_{n}),\ \tilde{\bm{u}}(t_{n})).$ (2.3)
@@
where $\tilde{\bm{u}}(t_{n})=\bm{u}_{t_{0}}+\int_{t_{0}}^{t_{n}}\tilde{\mathcal{G}}(\bm{u},t;\bm{\theta})\mathop{}\!\mathrm{d}t$ is the predicted states, $\bm{u}(t_{n})$ is the observed system states from the true system, and $J_{s}:\mathbb{R}^{d_{\bm{u}}\times M}\times\mathbb{R}^{d_{\bm{u}}\times M}\mapsto\mathbb{R}$ is a loss function at every time step. It is worth noting that the continuous counterpart of the total short-term loss $L_{s}$ is a loss functional operating on two space-time continuous functions and is defined by a squared norm for the Bochner space $L^{2}([0,T_{s}];L^{2}(D_{\bm{x}};\mathbb{R}^{d_{\bm{u}}}))$, where $T_{s}=t_{N_{s}}-t_{0}$ corresponds to the total time range of short-term trajectory matching. $N_{s}$ is a hyper-parameter controlling the short-term state prediction horizon during training process. Given a long time series of states as data of a dynamical system, we sample a batch of short time series with length $N_{s}$ from the dataset, with which we can use mini-batch training to obtain the parameters $\bm{\theta}$.
@@
To capture a long-term statistical property of the true dynamical system, the unknown parameters $\bm{\theta}$ in $\tilde{\mathcal{G}}$ can be calibrated by minimizing the long-term loss for statistics matching:
$L_{l}\left(\beta\left(\{\bm{u}\left(t_{n}\right)\}_{n=0}^{N_{l}}\right),\ \beta\left(\{\tilde{\bm{u}}\left(t_{n}\right)\}_{n=0}^{N_{l}}\right)\right).$ (2.4)
where $\beta:\mathbb{R}^{d_{\bm{u}}\times M\times N_{l}}\mapsto\mathbb{R}^{d_{\beta}}$ maps a long-term trajectory to its important statistics, $L_{l}:\mathbb{R}^{d_{\beta}}\times\mathbb{R}^{d_{\beta}}\mapsto\mathbb{R}$ is a loss function to quantify the differences of long-term statistics between the modeled system and the true one, and $N_{l}$ is another hyper-parameter that control the horizon over which the long-term statistics is calculated. The continuous counterpart of the map $\beta$ operates on a space-time continuous function in the space $C([0,T_{l}];L^{2}(D_{\bm{x}};\mathbb{R}^{d_{\bm{u}}}))$, where $T_{l}=t_{N_{l}}-t_{0}$ corresponds to the total time range of long-term statistics matching. In practice, to reduce the computational costs during the training process, we sample a batch of relatively long time series with length $N_{l}$ instead of calculating the long-term statistics for the entire time series dataset. The long-term horizon is typically much larger than the short-term horizon, i.e., $T_{l}\gg T_{s}$.